(b) Complete the Vertical Analysis table

Vertical Analysis

Income Statement

2018

2019

Gross Sales*

$950

$1,350

Cost of Goods Sold

$160

$175

Gross Margin

$790

$1,175

Expenses:

Salaries **

$150

$165

Warehousing Costs

$80

$80

Advertising ***

$60

$70

Bad Debt ****

$55

$55

Total expenses

$345

$370

Net Income

$445

$805

* Customer Returns amounted to $350 in each year

** Salaries include Commission and Bonus

*** Advertising includes newspaper, fliers, radio and TV ads

**** Allowances for Bad Debts typically represent 10% of Gross Sales
Answer #1

Vertical analysis will be prepared as follows

2018 2019
$ % $ %
Gross Sales* 950 158.33% 1350 135.00%
Less: Sales Returns -350 -58.33% -350 -35.00%
Net Sales 600 100.00% 1000 100.00%
Cost of Goods Sold 160 26.67% 175 17.50%
Gross Margin 790 131.67% 1175 117.50%
0.00%
Expenses: 0.00%
Salaries ** 150 25.00% 165 16.50%
Warehousing Costs 80 13.33% 80 8.00%
Advertising *** 60 10.00% 70 7.00%
Bad Debt **** 55 9.17% 55 5.50%
Total expenses 345 57.50% 370 37.00%
Net Income 445 74.17% 805 80.50%

Notes

To prepare vertical analysis of an income statement, all items of an income statement are expressed as a percentage of net revenues.

2018 2019 Gross Sales* Less: Sales Returns Net Sales Cost of Goods Sold Gross Margin 950 -350 =C95+C96 160 =C95-C98 =C95/$C$9

Please comment in case of any issue and I will be happy to help.